OpenStudy (anonymous):

Given: F = {(0, 1), (2, 4), (4, 6), (6, 8)} and G = {(2, 5), (4, 7), (5, 8), (6, 9), (7, 5)} (F + G) (2) = Can someone help me figure this out please!! Really Confused!!

OpenStudy (anonymous):

(f + g)(x) = f(x) + g(x)

OpenStudy (anonymous):

SO I can choose any ordered pair??

hartnn (hartnn):

no, find whats f(2) and g(2)

OpenStudy (anonymous):

ok thanks

hartnn (hartnn):

could u find f(2) ?

OpenStudy (anonymous):

whats the equation???

hartnn (hartnn):

no equation, you have to find it from the set

OpenStudy (anonymous):

how do you find it from the set?

hartnn (hartnn):

f(2) is the value at x=2 so search for that pair,which has 1st element = 2

OpenStudy (anonymous):

and....

hartnn (hartnn):

which pair did u get?

OpenStudy (anonymous):

did u get it as (2,4) for f(2)

OpenStudy (anonymous):

similarily find for g(2)

hartnn (hartnn):

@aussy123 which pair in f(x) did YOU get with 1st element as 2 ?

OpenStudy (anonymous):

(2,4)(2,5)

hartnn (hartnn):

BUT, (2,5) is in G isn't it?

OpenStudy (anonymous):

yea so (2,4) is chose (2,5) for G

hartnn (hartnn):

good, so f(2) will be 4 ok? and g(2) = 5 for (f+g)(2) just add them 4+5=?

OpenStudy (anonymous):

9 but is it just saying f(2)+g(2)=y; y1+y2=w/e

hartnn (hartnn):

whats w or e or w/e ?? 9 is the correct answer.

OpenStudy (anonymous):

whatever = w/e

OpenStudy (anonymous):

but is f(2)+g(2)=y; y1+y2=w/e The way to solve it

hartnn (hartnn):

ok :) y1=4,y2=5 y1+y2=4+5=9

OpenStudy (anonymous):

THANKS!!

hartnn (hartnn):

welcome :)